Doug lives for two periods. In the first period of his life he earns income Y1. The value of Y1 was determined by your student number. In the second period of his life, Doug is retired and does not earn any income. Doug’s decision is how much of his period one income should he save (S) in order to consume in period two. For every dollar that Doug saves in period one he has (1 + r) dollars available to spend in period two, where r is the interest rate.

Doug gets utility from consumption in period one (given by C1) and in period two (given by C2) according to the following utility function:
U(C1, C2) = ln(C1) + β ln(C2)

where the value of β was determined by your student number. Given this utility
function Doug’s marginal utility from consumption in period one is given by: MUC1 = ∂U / ∂C1 = 1 / C1

and his marginal utility from consumption in period two is given by:
MUC2 = ∂U / ∂C2 = β / C2

The parameter β describes how impatient Doug is. The lower the value of β the more he prefers consumption in the present (period one) to consumption when retired (period two).

(a) What is Doug’s budget constraint in period one? What is his budget constraint in period two?

(b) Combine these two budget constraints in order to have a single budget constraint that relates C1 and C2 to Doug’s income and the interest rate.

(c) Assume that Doug is currently saving exactly 40% of his income in period one and that the interest rate is 50%, that is r = 0.50.

i. How much does he consume in each period?
ii. Calculate the slope of Doug’s indifference curve at this point. Is the slope of his indifference curve at this point greater than, or less than the slope of his budget constraint?
iii. At this point could Doug raise his utility by increasing or decreasing the amount that he saves? Carefully explain your answer.

(d) Solve for Doug’s optimal choice of savings, and how much to consume in periods one and two.

Where:

β = 0.9
Y1= 20,000
Answer #1

u = lng + B In Ca al dn period. The can consume or tave. So G + s = 7, -0 Lie consumption & sowing a numel _ 0.is period 1 Bu[92=6= 0.box] slope I c = 1/2 slope of Rudget constraint = (1+r) =(1.50) Theres no way to figure out which has - greater sluThen Consumption decrease 4 to pay for in period 1 should full l. e .. 6 Increase as willingness G is less than the market to